The charge on the plates of a capacitor is 8mC when the potential
between them is 4 kV. Determine the capacitance of the capacitor.
1
Expert's answer
2018-11-28T11:17:11-0500

C=q/U=(8∙〖10〗^(-3))/(4∙〖10〗^3 )=2∙〖10〗^(-6) F=2 μF.